OpenStudy (anonymous):

3 11/12 < 1 2/4 ? 3 11/12 > 1 2/4 ? 3 11/12 = 1 2/4 ?

OpenStudy (hoblos):

3 11/12 = 47/12 1 2/4 = 6/4=18/12 ==>3 11/12 > 1 2/4

OpenStudy (anonymous):

I Need To Know Which Inequality Sign Is True ?

OpenStudy (hoblos):

yeah i wrote it it is 3 11/12 > 1 2/4

OpenStudy (anonymous):

Oh Ya Sorry Got Confused, Thanks (: